(a) At an air show a jet flies directly toward the stands at a speed of 1170 km/h, emitting a frequency of 3960 Hz, on a day when the speed of sound is 342 m/s. What frequency (in Hz) is received by the observers?

____________Hz

(b) What frequency (in Hz) do they receive as the plane flies directly away from them?

____________Hz
Answer #1

Solution) f = 3960 Hz

V = 1170 km/h

1 km/h = (1000)/(3600) m/s = (5/18) m/s

V = 1170×(5)/(18) = 325 m/s

Vs = 342 m/s

(a) f' = ?

f' = [f(Vs)]/(Vs - V)

f' = [3960×342]/(342 - 325)

f' = 79665.88 Hz

(b) f' = ? ( away)

f' = [f(Vs)]/(Vs + V)

f' = [3960×342]/(342+325)

f' = 2030.46 Hz